# From Local to Global: A Graph RAG Approach to Query-Focused Summarization

### v1

@@ Abstract
The use of retrieval-augmented generation (RAG) to retrieve relevant information from an external knowledge source enables large language models (LLMs) to answer questions over private and/or previously unseen document collections. However, RAG fails on global questions directed at an entire text corpus, such as “What are the main themes in the dataset?”, since this is inherently a query-focused summarization (QFS) task, rather than an explicit retrieval task. Prior QFS methods, meanwhile, fail to scale to the quantities of text indexed by typical RAG systems. To combine the strengths of these contrasting methods, we propose a Graph RAG approach to question answering over private text corpora that scales with both the generality of user questions and the quantity of source text to be indexed. Our approach uses an LLM to build a graph-based text index in two stages: first to derive an entity knowledge graph from the source documents, then to pregenerate community summaries for all groups of closely-related entities. Given a question, each community summary is used to generate a partial response, before all partial responses are again summarized in a final response to the user. For a class of global sensemaking questions over datasets in the 1 million token range, we show that Graph RAG leads to substantial improvements over a naïve RAG baseline for both the comprehensiveness and diversity of generated answers. An open-source, Python-based implementation of both global and local Graph RAG approaches is forthcoming at https://aka.ms/graphrag.
@@ 1 Introduction
Human endeavors across a range of domains rely on our ability to read and reason about large collections of documents, often reaching conclusions that go beyond anything stated in the source texts themselves. With the emergence of large language models (LLMs), we are already witnessing attempts to automate human-like sensemaking in complex domains like scientific discovery (Microsoft,, 2023) and intelligence analysis (Ranade and Joshi,, 2023), where sensemaking is defined as “a motivated, continuous effort to understand connections (which can be among people, places, and events) in order to anticipate their trajectories and act effectively” (Klein et al., 2006a, ). Supporting human-led sensemaking over entire text corpora, however, needs a way for people to both apply and refine their mental model of the data (Klein et al., 2006b, ) by asking questions of a global nature.
Retrieval-augmented generation (RAG, Lewis et al.,2020) is an established approach to answering user questions over entire datasets, but it is designed for situations where these answers are contained locally within regions of text whose retrieval provides sufficient grounding for the generation task. Instead, a more appropriate task framing is query-focused summarization (QFS, Dang,2006), and in particular, query-focused abstractive summarization that generates natural language summaries and not just concatenated excerpts (Yao et al.,, 2017; Baumel et al.,, 2018; Laskar et al.,, 2020) . In recent years, however, such distinctions between summarization tasks that are abstractive versus extractive, generic versus query-focused, and single-document versus multi-document, have become less relevant. While early applications of the transformer architecture showed substantial improvements on the state-of-the-art for all such summarization tasks (Liu and Lapata,, 2019; Laskar et al.,, 2022; Goodwin et al.,, 2020), these tasks are now trivialized by modern LLMs, including the GPT (Brown et al.,, 2020; Achiam et al.,, 2023), Llama (Touvron et al.,, 2023), and Gemini (Anil et al.,, 2023) series, all of which can use in-context learning to summarize any content provided in their context window.
Source Documents Text Chunks text extraction and chunking Element Instances domain-tailored summarization Element Summaries domain-tailored summarization Graph Communities community detection Community Summaries domain-tailored summarization Community Answers query-focused summarization Global Answer query-focused summarization Indexing Time Query Time Pipeline Stage
Figure 1: Graph RAG pipeline using an LLM-derived graph index of source document text. This index spans nodes (e.g., entities), edges (e.g., relationships), and covariates (e.g., claims) that have been detected, extracted, and summarized by LLM prompts tailored to the domain of the dataset. Community detection (e.g., Leiden, Traag et al.,, 2019) is used to partition the graph index into groups of elements (nodes, edges, covariates) that the LLM can summarize in parallel at both indexing time and query time. The “global answer” to a given query is produced using a final round of query-focused summarization over all community summaries reporting relevance to that query.
The challenge remains, however, for query-focused abstractive summarization over an entire corpus. Such volumes of text can greatly exceed the limits of LLM context windows, and the expansion of such windows may not be enough given that information can be “lost in the middle” of longer contexts (Liu et al.,, 2023; Kuratov et al.,, 2024). In addition, although the direct retrieval of text chunks in naïve RAG is likely inadequate for QFS tasks, it is possible that an alternative form of pre-indexing could support a new RAG approach specifically targeting global summarization.
In this paper, we present a Graph RAG approach based on global summarization of an LLM-derived knowledge graph (Figure 1). In contrast with related work that exploits the structured retrieval and traversal affordances of graph indexes (subsection 4.2), we focus on a previously unexplored quality of graphs in this context: their inherent modularity (Newman,, 2006) and the ability of community detection algorithms to partition graphs into modular communities of closely-related nodes (e.g., Louvain, Blondel et al.,2008; Leiden, Traag et al.,2019). LLM-generated summaries of these community descriptions provide complete coverage of the underlying graph index and the input documents it represents. Query-focused summarization of an entire corpus is then made possible using a map-reduce approach: first using each community summary to answer the query independently and in parallel, then summarizing all relevant partial answers into a final global answer.
To evaluate this approach, we used an LLM to generate a diverse set of activity-centered sensemaking questions from short descriptions of two representative real-world datasets, containing podcast transcripts and news articles respectively. For the target qualities of comprehensiveness, diversity, and empowerment (defined in subsection 3.4) that develop understanding of broad issues and themes, we both explore the impact of varying the the hierarchical level of community summaries used to answer queries, as well as compare to naïve RAG and global map-reduce summarization of source texts. We show that all global approaches outperform naïve RAG on comprehensiveness and diversity, and that Graph RAG with intermediate- and low-level community summaries shows favorable performance over source text summarization on these same metrics, at lower token costs.
2 Graph RAG Approach & Pipeline
@@
We now unpack the high-level data flow of the Graph RAG approach (Figure 1) and pipeline, describing key design parameters, techniques, and implementation details for each step.
2.1 Source Documents $\rightarrow$ Text Chunks
A fundamental design decision is the granularity with which input texts extracted from source documents should be split into text chunks for processing. In the following step, each of these chunks will be passed to a set of LLM prompts designed to extract the various elements of a graph index. Longer text chunks require fewer LLM calls for such extraction, but suffer from the recall degradation of longer LLM context windows (Liu et al.,, 2023; Kuratov et al.,, 2024). This behavior can be observed in Figure 2 in the case of a single extraction round (i.e., with zero gleanings): on a sample dataset (HotPotQA, Yang et al.,2018), using a chunk size of 600 token extracted almost twice as many entity references as when using a chunk size of 2400. While more references are generally better, any extraction process needs to balance recall and precision for the target activity.
01122330100002000030000Number of gleanings performedEntity references detected600 chunk size1200 chunk size2400 chunk size
varies with chunk size and gleanings for our generic entity extraction prompt with gpt-4-turbo.
2.2 Text Chunks $\rightarrow$ Element Instances
The baseline requirement for this step is to identify and extract instances of graph nodes and edges from each chunk of source text. We do this using a multipart LLM prompt that first identifies all entities in the text, including their name, type, and description, before identifying all relationships between clearly-related entities, including the source and target entities and a description of their relationship. Both kinds of element instance are output in a single list of delimited tuples.
@@ 2.2 Text Chunks $\rightarrow$ Element Instances
The primary opportunity to tailor this prompt to the domain of the document corpus lies in the choice of few-shot examples provided to the LLM for in-context learning (Brown et al.,, 2020). For example, while our default prompt extracting the broad class of “named entities” like people, places, and organizations is generally applicable, domains with specialized knowledge (e.g., science, medicine, law) will benefit from few-shot examples specialized to those domains. We also support a secondary extraction prompt for any additional covariates we would like to associate with the extracted node instances. Our default covariate prompt aims to extract claims linked to detected entities, including the subject, object, type, description, source text span, and start and end dates.
To balance the needs of efficiency and quality, we use multiple rounds of “gleanings”, up to a specified maximum, to encourage the LLM to detect any additional entities it may have missed on prior extraction rounds. This is a multi-stage process in which we first ask the LLM to assess whether all entities were extracted, using a logit bias of 100 to force a yes/no decision. If the LLM responds that entities were missed, then a continuation indicating that “MANY entities were missed in the last extraction” encourages the LLM to glean these missing entities. This approach allows us to use larger chunk sizes without a drop in quality (Figure 2) or the forced introduction of noise.
2.3 Element Instances $\rightarrow$ Element Summaries
The use of an LLM to “extract” descriptions of entities, relationships, and claims represented in source texts is already a form of abstractive summarization, relying on the LLM to create independently meaningful summaries of concepts that may be implied but not stated by the text itself (e.g., the presence of implied relationships). To convert all such instance-level summaries into single blocks of descriptive text for each graph element (i.e., entity node, relationship edge, and claim covariate) requires a further round of LLM summarization over matching groups of instances.
A potential concern at this stage is that the LLM may not consistently extract references to the same entity in the same text format, resulting in duplicate entity elements and thus duplicate nodes in the entity graph. However, since all closely-related “communities” of entities will be detected and summarized in the following step, and given that LLMs can understand the common entity behind multiple name variations, our overall approach is resilient to such variations provided there is sufficient connectivity from all variations to a shared set of closely-related entities.
Overall, our use of rich descriptive text for homogeneous nodes in a potentially noisy graph structure is aligned with both the capabilities of LLMs and the needs of global, query-focused summarization. These qualities also differentiate our graph index from typical knowledge graphs, which rely on concise and consistent knowledge triples (subject, predicate, object) for downstream reasoning tasks.
Refer to caption Refer to caption
(a) Root communities at level 0 (b) Sub-communities at level 1
2.4 Element Summaries $\rightarrow$ Graph Communities
The index created in the previous step can be modelled as an homogeneous undirected weighted graph in which entity nodes are connected by relationship edges, with edge weights representing the normalized counts of detected relationship instances. Given such a graph, a variety of community detection algorithms may be used to partition the graph into communities of nodes with stronger connections to one another than to the other nodes in the graph (e.g., see the surveys by Fortunato,2010 and Jin et al.,2021). In our pipeline, we use Leiden (Traag et al.,, 2019) on account of its ability to recover hierarchical community structure of large-scale graphs efficiently (Figure 3). Each level of this hierarchy provides a community partition that covers the nodes of the graph in a mutually-exclusive, collective-exhaustive way, enabling divide-and-conquer global summarization.
2.5 Graph Communities $\rightarrow$ Community Summaries

### v2

@@ Abstract
The use of retrieval-augmented generation (RAG) to retrieve relevant information from an external knowledge source enables large language models (LLMs) to answer questions over private and/or previously unseen document collections. However, RAG fails on global questions directed at an entire text corpus, such as “What are the main themes in the dataset?”, since this is inherently a query-focused summarization (QFS) task, rather than an explicit retrieval task. Prior QFS methods, meanwhile, do not scale to the quantities of text indexed by typical RAG systems. To combine the strengths of these contrasting methods, we propose GraphRAG, a graph-based approach to question answering over private text corpora that scales with both the generality of user questions and the quantity of source text. Our approach uses an LLM to build a graph index in two stages: first, to derive an entity knowledge graph from the source documents, then to pregenerate community summaries for all groups of closely related entities. Given a question, each community summary is used to generate a partial response, before all partial responses are again summarized in a final response to the user. For a class of global sensemaking questions over datasets in the 1 million token range, we show that GraphRAG leads to substantial improvements over a conventional RAG baseline for both the comprehensiveness and diversity of generated answers.
@@ 1 Introduction
Retrieval augmented generation (RAG) (Lewis et al.,, 2020) is an established approach to using LLMs to answer queries based on data that is too large to contain in a language model’s context window, meaning the maximum number of tokens (units of text) that can be processed by the LLM at once  (Liu et al.,, 2023; Kuratov et al.,, 2024). In the canonical RAG setup, the system has access to a large external corpus of text records and retrieves a subset of records that are individually relevant to the query and collectively small enough to fit into the context window of the LLM. The LLM then generates a response based on both the query and the retrieved records (Dang,, 2006; Yao et al.,, 2017; Baumel et al.,, 2018; Laskar et al.,, 2020). This conventional approach, which we collectively call vector RAG, works well for queries that can be answered with information localized within a small set of records. However, vector RAG approaches do not support sensemaking queries, meaning queries that require global understanding of the entire dataset, such as ”What are the key trends in how scientific discoveries are influenced by interdisciplinary research over the past decade?
Sensemaking tasks require reasoning over “connections (which can be among people, places, and events) in order to anticipate their trajectories and act effectively” (Klein et al.,, 2006). LLMs such as GPT (Brown et al.,, 2020; Achiam et al.,, 2023), Llama (Touvron et al.,, 2023), and Gemini (Anil et al.,, 2023) excel at sensemaking in complex domains like scientific discovery (Microsoft,, 2023) and intelligence analysis (Ranade and Joshi,, 2023). Given a sensemaking query and a text with an implicit and interconnected set of concepts, an LLM can generate a summary that answers the query. The challenge, however, arises when the volume of data requires a RAG approach, since vector RAG approaches are unable to support sensemaking over an entire corpus.
In this paper, we present GraphRAG – a graph-based RAG approach that enables sensemaking over the entirety of a large text corpus. GraphRAG first uses an LLM to construct a knowledge graph, where nodes correspond to key entities in the corpus and edges represent relationships between those entities. Next, it partitions the graph into a hierarchy of communities of closely related entities, before using an LLM to generate community-level summaries. These summaries are generated in a bottom-up manner following the hierarchical structure of extracted communities, with summaries at higher levels of the hierarchy recursively incorporating lower-level summaries. Together, these community summaries provide global descriptions and insights over the corpus. Finally, GraphRAG answers queries through map-reduce processing of community summaries; in the map step, the summaries are used to provide partial answers to the query independently and in parallel, then in the reduce step, the partial answers are combined and used to generate a final global answer.
The GraphRAG method and its ability to perform global sensemaking over an entire corpus form the main contribution of this work. To demonstrate this ability, we developed a novel application of the LLM-as-a-judge technique  (Zheng et al.,, 2024) suitable for questions targeting broad issues and themes where there is no ground-truth answer. This approach first uses one LLM to generate a diverse set of global sensemaking questions based on corpus-specific use cases, before using a second LLM to judge the answers of two different RAG systems using predefined criteria (defined in Section 3.3). We use this approach to compare GraphRAG to vector RAG on two representative real-world text datasets. Results show GraphRAG strongly outperforms vector RAG when using GPT-4 as the LLM.
GraphRAG is available as open-source software at https://github.com/microsoft/graphrag. In addition, versions of the GraphRAG approach are also available as extensions to multiple open-source libraries, including LangChain (LangChain,, 2024), LlamaIndex (LlamaIndex,, 2024), NebulaGraph (NebulaGraph,, 2024), and Neo4J (Neo4J,, 2024).
2 Background
2.1 RAG Approaches and Systems
RAG generally refers to any system where a user query is used to retrieve relevant information from external data sources, whereupon this information is incorporated into the generation of a response to the query by an LLM (or other generative AI model, such as a multi-media model). The query and retrieved records populate a prompt template, which is then passed to the LLM (Ram et al.,, 2023). RAG is ideal when the total number of records in a data source is too large to include in a single prompt to the LLM, i.e. the amount of text in the data source exceeds the LLM’s context window.
In canonical RAG approaches, the retrieval process returns a set number of records that are semantically similar to the query and the generated answer uses only the information in those retrieved records. A common approach to conventional RAG is to use text embeddings, retrieving records closest to the query in vector space where closeness corresponds to semantic similarity (Gao et al.,, 2023). While some RAG approaches may use alternative retrieval mechanisms, we collectively refer to the family of conventional approaches as vector RAG. GraphRAG contrasts with vector RAG in its ability to answer queries that require global sensemaking over the entire data corpus.
GraphRAG builds upon prior work on advanced RAG strategies. GraphRAG leverages summaries over large sections of the data source as a form of ”self-memory” (described in Cheng et al., 2024), which are later used to answer queries as in Mao et al., 2020). These summaries are generated in parallel and iteratively aggregated into global summaries, similar to prior techniques (Shao et al.,, 2023; Wang et al.,, 2024; Su et al.,, 2020; Feng et al.,, 2023; Trivedi et al.,, 2022; Khattab et al.,, 2022; Gao et al.,, 2023). In particular, GraphRAG is similar to other approaches that use hierarchical indexing to create summaries (similar to Kim et al., 2023; Sarthi et al., 2024). GraphRAG contrasts with these approaches by generating a graph index from the source data, then applying graph-based community detection to create a thematic partitioning of the data.
2.2 Using Knowledge Graphs with LLMs and RAG
Approaches to knowledge graph extraction from natural language text corpora include rule-matching, statistical pattern recognition, clustering, and embeddings (Yates et al.,, 2007; Mooney and Bunescu,, 2005; Kim et al.,, 2016; Etzioni et al.,, 2004). GraphRAG falls into a more recent body of research that use of LLMs for knowledge graph extraction (Melnyk et al.,, 2022; Tan et al.,, 2017; OpenAI,, 2023; Ban et al.,, 2023; Zhang et al., 2024a, ; Trajanoska et al.,, 2023; Yao et al.,, 2023; Yates et al.,, 2007). It also adds to a growing body of RAG approaches that use a knowledge graph as an index (Gao et al.,, 2023). Some techniques use subgraphs, elements of the graph, or properties of the graph structure directly in the prompt  (Baek et al.,, 2023; He et al.,, 2024; Zhang,, 2023) or as factual grounding for generated outputs (Kang et al.,, 2023; Ranade and Joshi,, 2023). Other techniques (Wang et al., 2023b, ) use the knowledge graph to enhance retrieval, where at query time an LLM-based agent dynamically traverses a graph with nodes representing document elements (e.g., passages, tables) and edges encoding lexical and semantical similarity or structural relationships. GraphRAG contrasts with these approaches by focusing on a previously unexplored quality of graphs in this context: their inherent modularity (Newman,, 2006) and the ability to partition graphs into nested modular communities of closely related nodes (e.g., Louvain, Blondel et al., 2008; Leiden, Traag et al., 2019). Specifically, GraphRAG recursively creates increasingly global summaries by using the LLM to create summaries spanning this community hierarchy.
2.3 Adaptive benchmarking for RAG Evaluation
Many benchmark datasets for open-domain question answering exist, including HotPotQA (Yang et al.,, 2018), MultiHop-RAG (Tang and Yang,, 2024), and MT-Bench (Zheng et al.,, 2024). However, these benchmarks are oriented towards vector RAG performance, i.e., they evaluate performance on explicit fact retrieval. In this work, we propose an approach for generating a set of questions for evaluating global sensemaking over the entirety of the corpus. Our approach is related to LLM methods that use a corpus to generate questions whose answers would be summaries of the corpus, such as in Xu and Lapata, (2021). However, in order to produce a fair evaluation, our method avoids generating the questions directly from the corpus itself (as an alternative implementation, one can use a subset of the corpus held out from subsequent graph extraction and answer evaluation steps).
Adaptive benchmarking refers to the process of dynamically generating evaluation benchmarks tailored to specific domains or use cases. Recent work has used LLMs for adaptive benchmarking to ensure relevance, diversity, and alignment with the target application or task (Yuan et al.,, 2024; Zhang et al., 2024b, ). In this work, we propose an adaptive benchmarking approach to generating global sensemaking queries for the LLM. Our approach builds on prior work in LLM-based persona generation, where the LLM is used to generate diverse and authentic sets of personas (Salminen et al.,, 2024; Shin et al.,, 2024; Kosinski,, 2024). Our adaptive benchmarking procedure uses persona generation to create queries that are representative of real-world RAG system usage. Specifically, our approach uses the LLM to infer the potential users would use the RAG system and their use cases, which guide the generation of corpus-specific sensemaking queries.
2.4 RAG evaluation criteria
Our evaluation relies on the LLM to evaluate how well the RAG system answers the generated questions. Prior work has shown LLMs to be good evaluators of natural language generation, including work where LLMs evaluations were competitive with human evaluations (Wang et al., 2023a, ; Zheng et al.,, 2024). Some prior work proposes criteria for having LLMs quantify the quality of generated texts such as “fluency”  (Wang et al., 2023a, ) Some of these criteria are generic to vector RAG systems and not relevant to global sensemaking, such as “context relevance”, “faithfulness”, and “answer relevance” (RAGAS, Es et al., 2023). Lacking a gold standard for evaluation, one can quantify relative performance for a given criterion by prompting the LLM to compare generations from two different competing models (LLM-as-a-judge, (Zheng et al.,, 2024)). In this work, we design criteria for evaluating RAG-generated answers to global sensemaking questions and evaluate our results using the comparative approach. We also validate results using statistics derived from LLM-extracted statements of verifiable facts, or “claims.”
3 Methods
3.1 GraphRAG Workflow
@@
Figure 1 illustrates the high-level data flow of the GraphRAG approach and pipeline. In this section, we describe the key design parameters, techniques, and implementation details for each step.
Source Documents Text Chunks text extraction and chunking Entities & Relationships domain-tailored summarization Knowledge Graph domain-tailored summarization Graph Communities community detection Community Summaries domain-tailored summarization Community Answers query-focused summarization Global Answer query-focused summarization Indexing Time Query Time Pipeline Stage
Figure 1: Graph RAG pipeline using an LLM-derived graph index of source document text. This graph index spans nodes (e.g., entities), edges (e.g., relationships), and covariates (e.g., claims) that have been detected, extracted, and summarized by LLM prompts tailored to the domain of the dataset. Community detection (e.g., Leiden, Traag et al.,, 2019) is used to partition the graph index into groups of elements (nodes, edges, covariates) that the LLM can summarize in parallel at both indexing time and query time. The “global answer” to a given query is produced using a final round of query-focused summarization over all community summaries reporting relevance to that query.
3.1.1 Source Documents $\rightarrow$ Text Chunks
To start, the documents in the corpus are split into text chunks. The LLM extracts information from each chunk for downstream processing. Selecting the size of the chunk is a fundamental design decision; longer text chunks require fewer LLM calls for such extraction (which reduces cost) but suffer from degraded recall of information that appears early in the chunk (Liu et al.,, 2023; Kuratov et al.,, 2024). See Section A.1 for prompts and examples of the recall-precision trade-offs.
3.1.2 Text Chunks $\rightarrow$ Entities & Relationships
In this step, the LLM is prompted to extract instances of important entities and the relationships between the entities from a given chunk. Additionally, the LLM generates short descriptions for the entities and relationships. To illustrate, suppose a chunk contained the following text:
NeoChip’s (NC) shares surged in their first week of trading on the NewTech Exchange. However, market analysts caution that the chipmaker’s public debut may not reflect trends for other technology IPOs. NeoChip, previously a private entity, was acquired by Quantum Systems in 2016. The innovative semiconductor firm specializes in low-power processors for wearables and IoT devices.
The LLM is prompted such that it extracts the following:
  • The entity NeoChip, with description “NeoChip is a publicly traded company specializing in low-power processors for wearables and IoT devices.”
  • The entity Quantum Systems, with description “Quantum Systems is a firm that previously owned NeoChip.”
  • A relationship between NeoChip and Quantum Systems, with description “Quantum Systems owned NeoChip from 2016 until NeoChip became publicly traded.”
@@ 3.1.2 Text Chunks $\rightarrow$ Entities & Relationships
These prompts can be tailored to the domain of the document corpus by choosing domain appropriate few-shot exemplars for in-context learning (Brown et al.,, 2020). For example, while our default prompt extracts the broad class of “named entities” like people, places, and organizations and is generally applicable, domains with specialized knowledge (e.g., science, medicine, law) will benefit from few-shot exemplars specialized to those domains.
The LLM can also be prompted to extract claims about detected entities. Claims are important factual statements about entities, such as dates, events, and interactions with other entities. As with entities and relationships, in-context learning exemplars can provide domain-specific guidance. Claim descriptions extracted from the example tetx chunk are as follows:
  • NeoChip’s shares surged during their first week of trading on the NewTech Exchange.
  • NeoChip debuted as a publicly listed company on the NewTech Exchange.
  • Quantum Systems acquired NeoChip in 2016 and held ownership until NeoChip went public.
See Appendix A for prompts and details on our implementation of entity and claim extraction.
3.1.3 Entities & Relationships $\rightarrow$ Knowledge Graph
The use of an LLM to extract entities, relationships, and claims is a form of abstractive summarization – these are meaningful summaries of concepts that, in the case of relationships and claims, may not be explicitly stated in the text. The entity/relationship/claim extraction processes creates multiple instances of a single element because an element is typically detected and extracted multiple times across documents.
In the final step of the knowledge graph extraction process, these instances of entities and relationships become individual nodes and edges in the graph. Entity descriptions are aggregated and summarized for each node and edge. Relationships are aggregated into graph edges, where the number of duplicates for a given relationship becomes edge weights. Claims are aggregated similarly.
In this manuscript, our analysis uses exact string matching for entity matching – the task of reconciling different extracted names for the same entity (Barlaug and Gulla,, 2021; Christen and Christen,, 2012; Elmagarmid et al.,, 2006). However, softer matching approaches can be used with minor adjustments to prompts or code. Furthermore, GraphRAG is generally resilient to duplicate entities since duplicates are typically clustered together for summarization in subsequent steps.
3.1.4 Knowledge Graph $\rightarrow$ Graph Communities
Given the graph index created in the previous step, a variety of community detection algorithms may be used to partition the graph into communities of strongly connected nodes (e.g., see the surveys by Fortunato, (2010) and Jin et al., (2021)). In our pipeline, we use Leiden community detection (Traag et al.,, 2019) in a hierarchical manner, recursively detecting sub-communities within each detected community until reaching leaf communities that can no longer be partitioned.
Each level of this hierarchy provides a community partition that covers the nodes of the graph in a mutually exclusive, collectively exhaustive way, enabling divide-and-conquer global summarization. An illustration of such hierarchical partitioning on an example dataset can be found in Appendix B.
3.1.5 Graph Communities $\rightarrow$ Community Summaries